Show that the sequence {fn}of functions, where

fn (x )=n/(x+2n)
is uniformly
convergent in [0.k] where k > 0.
1
Expert's answer
2020-03-04T18:12:30-0500

Definition. If we define


"d_n=\sup\limits_{x\in E}|f_n(x)-f(x)|"


then "f_n" converges to "f"  uniformly if and only if "{\displaystyle d_{n}\to 0}"  as "{\displaystyle n\to \infty }" .

( More information: https://en.wikipedia.org/wiki/Uniform_convergence )

In our case,

1 STEP:


"\lim\limits_{n\to\infty}f_n(x)=\lim\limits_{n\to\infty}\left(\frac{n}{x+2n}\right)=\frac{1}{2}"

2 STEP:


"\left|\frac{n}{x+2n}-\frac{1}{2}\right|=\left|\frac{2n-(x+2n)}{2(x+2n)}\right|=\left|\frac{-x}{2(x+2n)}\right|=\\[0.5cm]
\frac{x}{2(x+2n)}"

3 STEP: we must find "sup" for the expression above on the interval "x\in[0;k], \forall k>0".

To do this, we will look at the expression above as a certain function "y(x)" and use the derivative.


"y'(x)=\frac{d}{dx}\left(\frac{x}{2(x+2n)}\right)=\frac{1\cdot(2(x+2n))-x\cdot2}{(2(x+2n))^2}=\\[0.5cm]
=\frac{4n}{4(x+2n)^2}=\frac{n}{(x+2n)^2}>0, \forall x\in[0;k],\forall n\in\mathbb{N}."

Then,


"d_n=\sup\limits_{x\in[0;k]}y(x)=\sup\limits_{x\in[0;k]}\frac{x}{2(x+2n)}=\frac{k}{2(k+2n)}"

3 STEP:


"\lim\limits_{n\to\infty}d_n=\lim\limits_{n\to\infty}\frac{k}{2(k+2n)}=0"

Conclusion,


"\boxed{f_n(x)=\frac{x}{x+2n}\rightrightarrows\frac{1}{2}}"

(the sequence is uniformly convergent).